OpenStudy (anonymous):

AD/AB=AB/AC. If AD=9 cm and DC=4 cm, then AB= ___ and BC= ___. Anybody know the answer?

OpenStudy (anonymous):

find AC \[9^2+4^2=97\] \[\sqrt{97}=9.8\] 9.8*9=88.2 88.2=AB*AB \[AB=\sqrt{88.2}=9.4\] \[\sqrt{9.8^2+9.4^2}=BC\] =2.8